OpenStudy (anonymous):

(3x over x-5) - (5 over 5-x)

OpenStudy (anonymous):

\[{3x \over x-5} - {5 \over 5-x} \]

OpenStudy (anonymous):

yea

OpenStudy (ghazi):

\[\frac{ 3x }{ x-5 }- \frac{ 5 }{ 5-x }= \frac{ 3x }{ x-5 }+ \frac{ 5 }{ x-5 }\]

OpenStudy (anonymous):

its the same thing?

OpenStudy (ghazi):

\[\frac{ 3x+5 }{ x-5 }\]

OpenStudy (anonymous):

okay so you change it so 3x+5 to change the 5-x to x-5 right?

OpenStudy (anonymous):

He distributed the negative into the second equation to make the negatives positive which resulted in the same denominator: x-5. (I believe that is what you're asking?)

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Ok.